find two conseutive odd possitive odd no. whish are such that the square of their sum exceeds the sum of their square by 126
Answer by mananth(16949)

let the numbers be x & x+2
the sum of their squares = x^2 + (x+2)^2
the square of their sum = (x+x+2)^2
(2x+2)^2=x^2+(x+2)^2+126
4x^2+8x+4=x^2+x^2+4x+4 +126
2x^2+4x-126=0
x^2+2x-63=0
x^2+9x-7x-63=0
x(x+9)-7(x+9)=0
(x-7)(x+9)=0
x=7 OR x=-9
7 & 9